# Spherically symmetric solutions to the Einstein-scalar field conformal constraint equations

### v1

@@ 4.4. Stability and Instability of conformal equations
Theorem 4.11 (Positive $\mathcal{B}_{\tau,\psi}$ and stability).
@@
  • for each $i$, $(\varphi_{i}(r),W_{i})$ is solution to the conformal constraint equations with data set $V=\Lambda$, $\tau=\tau_{i}$, $\psi\equiv 0$ and $\rho\equiv 0$;
@@
$\tfrac{2n\Lambda}{n-1}b^{N-2}>(\sin^{a}(r)+\varepsilon)^{N-2-\frac{2}{a}}\bigg((\sin^{a}(r)+\varepsilon)^{{2\over a}}-\frac{N^{2}a^{2}(\sin^{a}(r))^{2-\frac{2}{a}}\cos^{2}(r)}{(\sin^{a}(r)+\varepsilon)^{2-\frac{2}{a}}}\\ +\frac{2nNa(\sin^{a}(r))^{1-\frac{2}{a}}\cos^{2}(r)}{(\sin^{a}(r)+\varepsilon)^{1-\frac{2}{a}}}+\frac{n^{2}}{(\sin^{a}(r)+\varepsilon)^{-\frac{2}{a}}}\bigg).$ (4.54)
@@ 5. Solutions in the hyperbolic space
Building on this approach, besides proving Main Theorem 1 in hyperbolic spaces, which provides a necessary and sufficient condition for a positive function $\varphi$ to solve (3.13), an important outcome in this section is that, as long as $\tau$ does not change sign, the conformal equations in radial settings on hyperbolic spaces are shown to be solvable, with the set of solutions being compact. Although this result is limited to radial settings, it provides strong evidence for the solvability of (1.2) in general data, with the limit equation criterion offering a promising tool. This indicates that, in contrast to the conclusions obtained in [43, 44, 18] for compact manifolds, the conformal method remains a powerful approach for parametrizing constraint solutions on AH spaces, not only in the CMC or near-CMC regimes, but also in the far-from-CMC regime.
@@
The organization of the section is as follows. In Subsection 5.1, we recall some basic notation on hyperbolic space and then provide the proof of Main Theorem 1 in this setting. In Subsection 5.2, provided that $\tau$ does not change sign, the solvability of the conformal equations in radial data settings will be obtained. Finally, in the last subsection, we review the conjecture on the positivity of the AH mass and show how the decay rate of constraint solutions at infinity derives the sign of the mass.
@@ 5.1. Lichnerowicz-type equation in the hyperbolic space
    If $S_{V,\psi,\rho,\varphi}>0$ on $\mathbb{R}_{+}$ and $(\varphi^{N}\sinh^{n})^{\prime}\neq 0$ a.e. on $\mathbb{R}_{+}$, then $\varphi$ is solution to (5.2) if and only if the function $\tau$ satisfies
@@
The proof of this theorem is broadly similar to that of Theorem 4.1 line by line, with the only difference being that we replace $(\mathrm{s}_{\kappa},\mathrm{c}_{\kappa})=(\sin(r),\cos(r))$ on the sphere by $(\sinh(r),\cosh(r))$ in hyperbolic space, as defined in (1.5). Here, for the reader’s convenience, we present the proof as follows.
@@
Similarly to Appendix A, it is not difficult to find a set of $(V,\psi,\rho,\varphi)$ such that $S_{V,\psi,\rho,\varphi}$ is positive on hyperbolic spaces. Therefore, thanks to the theorem, a large class of explicit solutions to (1.2) can be easily obtained, and so it gives a variety of models in general relativity, which can help us to find out important properties of initial data in both theoretical and numerical sense. In particular, in the vacuum case, the condition $S_{V,\psi,\rho,\varphi}>0$ can be considerably simplified, making it quite simple to choose $\varphi$ so that the condition is satisfied. Since this result will be used later in the discussion of the sign of the AH mass, we write it here as an observation.
@@
If $V\equiv 0$ and $\psi=\rho=0$, then $S_{V,\psi,\rho,\varphi}>0$ if and only if $\big(\big(\cosh({r\over 2})\big)^{n-2}\varphi\big)^{\prime}>0$. In fact, in this case, by definition, $S_{V,\psi,\rho,\varphi}>0$ is rewritten as
@@
$\bigg({N\over n}{\varphi^{\prime}\over\varphi}\bigg)^{2}+2\coth(r)\bigg({N\over n}{\varphi^{\prime}\over\varphi}\bigg)+1>0.$ (5.10)
@@
${N\over n}{\varphi^{\prime}\over\varphi}>-\tanh\bigg({r\over 2}\bigg).$
By straightforward calculations, this is equivalent to $\bigg(\big(\cosh({r\over 2})\big)^{n-2}\varphi\bigg)^{\prime}>0$, which is our claim.
@@
$\varphi:=1\pm{1\over kq}\bigg({r^{q}\over r^{q}+1}\bigg)e^{-pr},$ (5.11)
@@ 5.2. Existence of solutions with mean curvature of constant sign
It will be clear from the context if the notation refers to a space of functions on $\mathbb{H}^{n}$, or a space of sections of some bundle over $\mathbb{H}^{n}$. We can easily check that $(C^{l,\alpha}_{-p},\|.\|_{C^{l,\alpha}_{-p}})$ satisfies the weighted Sobolev embedding and Rellich theorem on $\mathbb{H}^{n}$. A reference for this can be found in [25, Proposition 2.3].
@@
with $\varphi\longrightarrow 1$ at infinity. Heuristically, the analysis of (5.13) is essentially the same as that of the vacuum system (1.2) studied in [31, 24, 25], where the existence of solutions is known only when $\tau$ is constant or near constant. Therefore, to obtain a more global understanding of the equations—particularly to seek convincing evidence as to whether (1.2) remains solvable in the far-from-CMC regime— we will consider (5.13) in the simpler setting where both $\tau$ and $\rho$ are radial. In this framework, since the analysis takes place within the space of radial functions as we have seen in Theorem 5.1, rather than seeking a general solution $\varphi$ to (5.13), it is natural to restrict our consideration to the set of radial solutions.
In this context, thanks to the limit equation criterion established by R. Gicquaud and A. Sakovich [25], we will follow the approach of Theorem 4.9 and show that, as long as $\tau$ does not change sign, the conformal equations (5.13) are always solvable, with expect properties of solutions satisfied. More precisely, the statement of our result is as follows.
@@
Then, the conformal equations (5.13) admit at least one solution $\varphi(r)>0$ with $\varphi(r)-1\in C^{2,\alpha}_{-p}(\mathbb{H}^{n})$. Moreover, in this case, the set of radial solutions is compact.
@@
Following R. Gicquaud and A. Sakovich [25, Section 3.3], we obtain that $T$ is continuous and compact. It is clear that a fixed point of $T$ is a solution to (5.13), therefore, the theorem will follows if we can show that $T$ has a fixed point. We will argue by contradiction. In fact, assume that $T$ has no fixed point. Then, according to the Leray–Schauder’s fixed point theorem, there exists a sequence $\{(t_{i},\phi_{i}(r))\}\subset(0,1)\times RC^{2,\alpha}_{-p}(\mathbb{H}^{n})$ such that
@@
In view of [25, Section 4], when $i\to+\infty$, similarly to the proof of Theorem 4.9, (up to a subsequence) there exists $y(r)\in RC^{0}(\mathbb{H}^{n})\setminus\{0\}$ and $t_{0}\in(0,1]$ such that, when $i\to+\infty$,
@@
5.3. Sharpness of the decay rate in the Positive Mass Theorem
@@
Thanks to Theorem 5.1 and the conformal method, it is clear that $(g,k)$ defined in (5.20) is a solution to the vacuum constraint (1.1) on the hyperbolic manifold. Let us now consider the mass of $(g,k)$. We first compute $e$ and $\pi$ from their definitions
@@
$\lim\limits_{r\to+\infty}a^{\prime}(r)=\lim\limits_{r\to+\infty}{a\over r}=0$
@@
where $\ell=\lim\limits_{r\to+\infty}a(r)$. Here, we observe that, as can be easily verified through Example (5.11) in Remark 5.2, this limit $\ell$ can have an arbitrary sign. Therefore, it follows in particular that if $p=\frac{n}{2}$, then $\mathcal{M}_{(g,k)}$ is finite and can also have an arbitrary sign, and if $p<\frac{n}{2}$, then $\mathcal{M}_{(g,k)}$ can reach $\pm\infty$ as claimed. The proof is completed. ∎
@@
It is worth noting that since the vacuum solution $(g,k)$ constructed in the proof of Theorem 5.4 is spherically symmetric and regular in $\mathbb{R}^{+}$, by Birkhoff’s theorem, such a $(\mathbb{H}^{n},g,k)$ in the theorem is indeed Cauchy initial data for Minkowski space.
@@ 6.1. Lichnerowicz-type equation in the Euclidean space
In the same manner as in the previous sections, from a radial data set $(V,\tau(r),\psi(r),\rho(r))\in C^{0}(\mathbb{R})\times RC^{1}(\mathbb{R}^{n})\times RC^{1}(\mathbb{R}^{n})\times RC^{0}(\mathbb{R}^{n})$ and a radial function $\varphi(r)\in RC_{+}^{3}(\mathbb{R}^{n})$ we define:
@@
    If $S_{V,\psi,\rho,\varphi}>0$ on $\mathbb{R}_{+}$ and $(\varphi^{N}r^{n})^{\prime}\neq 0$ a.e on $\mathbb{R}_{+}$, then $\varphi$ is solution to (6.2) if and only if the function $\tau$ satisfies
@@ 6.2. Existence of solutions with freely specified mean curvature
Although the Euclidean space is the most familiar setting, existence results for (1.2) are almost restricted to the CMC case. This is because, apart from the limit equation criterion—which, as explained at the beginning of this section, has not yet been established for the conformal equations on AF manifolds—there are essentially only two approaches available for proving existence of solutions in the non-CMC regime. The first relies on the smallness of $d\tau\over\tau$} to prevent blow-up of solutions, corresponding to the near-CMC case. The second is based on the smallness of $\sigma$ and $\rho$, leading to small solutions and is commonly called the small TT-tensor method. Unfortunately, despite the results in [17], none of these methods is convincingly applicable on $\mathbb{R}^{n}$. On the one hand, the weighted Poincaré inequality does not allow us to assume that $\frac{d\tau}{\tau}$ is too small; on the other hand, the fact that solutions converges to $1$ at infinity breaks the smallness assumption on solutions required to implement the small TT-tensor method.
@@
with $\varphi\longrightarrow 1$ at infinity. The close analytical similarity of these equations to the vacuum (1.2) and their importance in providing a global perspective to the solvability of (1.2) have been explained briefly in the previous sections. For that, similarly to what we have done previously, to establish convincing evidence supporting that the conformal method is still an useful tool for constructing solutions to the constraint (1.1) on AF manifolds, we study (6.4) particularly in the simple model where $(M,g)$ is Euclidean and $(\tau,\rho)$ are assumed to be radial. In this context, motivated by the proof of Theorem 4.10, the main result we obtain in this section is that as long as $\rho$ is $C^{1}$-regular, the equations (6.4) are always solvable, with solutions uniformly bounded independently of $\tau$. Before presenting the precis statement and providing the proof, we first introduce the associated Banach spaces on $(\mathbb{R}^{n},\sigma_{\text{Euc}})$ as follows. Given an integer $l\geq 0$, a Hölder exponent $\alpha\in[0,1]$, and a decay exponent $\beta>0$, we will use the weighted Hölder spaces $C^{l,\alpha}_{-\beta}$ to capture asymptotic of functions and tensors near infinity. For $\alpha=0$, we will write $C^{l}_{-\beta}$ instead of $C^{l,0}_{-\beta}$. The weighted norm convention we are using is that the $C^{s,\alpha}_{-\beta}$ norm is given by
@@
In view of [16, Section 5], we obtain that $\widetilde{T}$ is continuous compact and it is clear that a fixed point of $\widetilde{T}(1,)$ is a solution to (6.4). Now, we set
@@
${\rho^{2}(\varphi^{N}r^{n})^{\prime}\over\varphi^{2N}}=-(\rho^{2}r^{2n})\bigg({1\over\varphi^{N}r^{n}}\bigg)^{\prime},$
@@
6.3. Sharpness of the decay rate in the Positive Mass Theorem

### v2

@@ 4.4. Stability and Instability of conformal equations
Theorem 4.11 (Non-positive $\mathcal{B}_{\tau,\psi}$ and stability).
@@
  • for each $i$, $(\varphi_{i}(r),W_{i})$ is a solution to the conformal constraint equations with data set $V=\Lambda$, $\tau=\tau_{i}$, $\psi\equiv 0$ and $\rho\equiv 0$;
@@
$\tfrac{2n\Lambda}{n-1}b^{N-2}>(\sin^{a}(r)+\varepsilon)^{N-2-\frac{2}{a}}\bigg((\sin^{a}(r)+\varepsilon)^{{\frac{2}{a}}}-\frac{N^{2}a^{2}(\sin^{a}(r))^{2-\frac{2}{a}}\cos^{2}(r)}{(\sin^{a}(r)+\varepsilon)^{2-\frac{2}{a}}}\\ +\frac{2nNa(\sin^{a}(r))^{1-\frac{2}{a}}\cos^{2}(r)}{(\sin^{a}(r)+\varepsilon)^{1-\frac{2}{a}}}+\frac{n^{2}}{(\sin^{a}(r)+\varepsilon)^{-\frac{2}{a}}}\bigg).$ (4.54)
@@ 5. Solutions in the hyperbolic space
Building on this approach, besides proving Main Theorem 1 in hyperbolic spaces, which provides a necessary and sufficient condition for a positive function $\varphi$ to solve (3.13), an important outcome in this section is that, as long as $\tau$ does not change sign, the conformal equations in the radial setting on hyperbolic space are shown to be solvable, with a compact set of solutions. Although this result is limited to the radial setting, it provides strong evidence for the solvability of (1.2) in general data, with the limit equation criterion offering a promising tool. This indicates that, in contrast to the conclusions obtained in [36, 37, 15] for compact manifolds, the conformal method remains a powerful approach for parametrizing constraint solutions on AH spaces, not only in the CMC or near-CMC regimes, but also in the far-from-CMC regime.
@@
The organization of the section is as follows. In Subsection 5.1, we recall some basic notation on hyperbolic space and then provide the proof of Main Theorem 1 in this setting. In Subsection 5.2, provided that $\tau$ does not change sign, the solvability of the conformal equations in the radial data setting will be obtained. Finally, in the last subsection, we review the conjecture on the positivity of the AH mass and show how the decay rate of constraint solutions at infinity derives the sign of the mass.
@@ 5.1. Lichnerowicz-type equation in the hyperbolic space
    If $S_{V,\psi,\rho,\varphi}>0$ on $\mathbb{R}_{+}$ and $(\varphi^{N}\sinh^{n})^{\prime}\neq 0$ a.e. on $\mathbb{R}_{+}$, then $\varphi$ is a solution to (5.2) if and only if the function $\tau$ satisfies
@@
The proof of this theorem is broadly similar to that of Theorem 4.1 line by line, with the only difference being that we replace $(\mathrm{s}_{\kappa},\mathrm{c}_{\kappa})=(\sin(r),\cos(r))$ on the sphere by $(\sinh(r),\cosh(r))$ in hyperbolic space, as defined in (1.5). Here, for the reader’s convenience, we detail the proof as follows.
@@
Similarly to Appendix A, it is not difficult to find a set of $(V,\psi,\rho,\varphi)$ such that $S_{V,\psi,\rho,\varphi}$ is positive on hyperbolic spaces. Therefore, thanks to the theorem, a large class of explicit solutions to (1.2) can be easily obtained, and so it gives a variety of models in general relativity, which can help us to find out important properties of initial data in both theoretical and numerical sense.
@@
In the vacuum case, the condition $S_{V,\psi,\rho,\varphi}>0$ can be considerably simplified, making it quite simple to choose $\varphi$ so that the condition is satisfied. Since this result will be used later in the discussion of the sign of the AH mass, we write it here as an observation.
If $V\equiv 0$ and $\psi=\rho=0$, then $S_{V,\psi,\rho,\varphi}>0$ if and only if $\big(\big(\cosh(\frac{r}{2})\big)^{n-2}\varphi\big)^{\prime}>0$. In fact, in this case, by definition, $S_{V,\psi,\rho,\varphi}>0$ is rewritten as
@@
$\bigg(\tfrac{N}{n}\frac{\varphi^{\prime}}{\varphi}\bigg)^{2}+2\coth(r)\bigg(\tfrac{N}{n}\frac{\varphi^{\prime}}{\varphi}\bigg)+1>0.$ (5.10)
@@
$\tfrac{N}{n}\frac{\varphi^{\prime}}{\varphi}>-\tanh\big(\tfrac{r}{2}\big).$
By straightforward calculations, this is equivalent to $\bigg(\big(\cosh(\tfrac{r}{2})\big)^{n-2}\varphi\bigg)^{\prime}>0$, which is our claim.
@@
$\varphi:=1\pm\frac{1}{kq}\bigg(\frac{r^{q}}{r^{q}+1}\bigg)e^{-pr},$ (5.11)
@@ 5.2. Existence of solutions with mean curvature of constant sign
It will be clear from the context if the notation refers to a space of functions on $\mathbb{H}^{n}$, or a space of sections of some bundle over $\mathbb{H}^{n}$. We can easily check that $(C^{l,\alpha}_{-p},\|.\|_{C^{l,\alpha}_{-p}})$ satisfies the weighted Sobolev embedding and Rellich theorem on $\mathbb{H}^{n}$. A reference for this can be found in [21, Proposition 2.3].
@@
with $\varphi\longrightarrow 1$ at infinity. Heuristically, the analysis of (5.13) is essentially the same as that of the vacuum system (1.2) studied in [26, 20, 21], where the existence of solutions is known only when $\tau$ is constant or near constant. Therefore, to obtain a more global understanding of the equations — particularly to seek convincing evidence to the existence of solutions to (1.2) in the far-from-CMC regime — we will consider (5.13) in the simpler setting where both $\tau$ and $\rho$ are radial. In this framework, since the analysis takes place within the space of radial functions as we have seen in Theorem 5.1, rather than seeking a general solution $\varphi$ to (5.13), it is natural to restrict our consideration to the set of radial solutions.
In this context, thanks to the limit equation criterion established by R. Gicquaud and A. Sakovich [21], we will follow the approach of Theorem 4.9 and show that, as long as $\tau$ does not change sign, the conformal equations (5.13) are always solvable, with solutions satisfying the expected properties. More precisely, the statement of our result is as follows.
@@
Then, the conformal equations (5.13) admit at least one solution $\varphi(r)>0$ with $\varphi(r)-1\in C^{2,\alpha}_{-p}(\mathbb{H}^{n})$. Moreover, in this case, the set of solutions is compact.
@@
Following R. Gicquaud and A. Sakovich [21, Section 3.3], we obtain that $T$ is continuous and compact. It is clear that a fixed point of $T$ is a solution to (5.13), therefore, the theorem will follows if we can show that $T$ has a fixed point. We will argue by contradiction. In fact, assume that $T$ has no fixed point. Then, according to the Leray–Schauder’s fixed point theorem, there exists a sequence $\{(t_{i},\phi_{i}(r))\}\subset(0,1)\times RC^{2,\alpha}_{-p}(\mathbb{H}^{n})$ such that
@@
In view of [21, Section 4], when $i\to+\infty$, similarly to the proof of Theorem 4.9, (up to a subsequence) there exists $y(r)\in RC^{0}(\mathbb{H}^{n})\setminus\{0\}$ and $t_{0}\in(0,1]$ such that, when $i\to+\infty$,
@@
5.3. The sharpness of the decay-exponent in the positive mass theorem
@@
Thanks to Theorem 5.1 and the conformal method, it is clear that $(g,k)$ defined in (5.20) is a solution to the vacuum constraint equations (1.1) on the hyperbolic manifold. Let us now consider the mass of $(g,k)$. We first compute $e$ and $\pi$ from their definitions
@@
$\lim\limits_{r\to+\infty}a^{\prime}(r)=\lim\limits_{r\to+\infty}\frac{a}{r}=0$
@@
where $\ell=\lim\limits_{r\to+\infty}a(r)$. Here, we observe that, as can be easily verified through Example 5.11 in Remark 5.2, this limit $\ell$ can have an arbitrary sign. Therefore, it follows in particular that if $p=\frac{n}{2}$, then $\mathcal{M}_{(g,k)}$ is finite and can also have an arbitrary sign, and if $p<\frac{n}{2}$, then $\mathcal{M}_{(g,k)}$ can reach $\pm\infty$ as claimed. The proof is completed. ∎
@@
It is worth noting that since the vacuum solution $(g,k)$ constructed in the proof of Theorem 5.4 is spherically symmetric and regular, by Birkhoff’s theorem, such a $(\mathbb{H}^{n},g,k)$ in the theorem is indeed Cauchy initial data for Minkowski space.
@@ 6.1. Lichnerowicz-type equation in the Euclidean space
As in the previous sections, from a radial data set $(V,\tau(r),\psi(r),\rho(r))\in C^{0}(\mathbb{R})\times RC^{1}(\mathbb{R}^{n})\times RC^{1}(\mathbb{R}^{n})\times RC^{0}(\mathbb{R}^{n})$ and a radial function $\varphi(r)\in RC_{+}^{3}(\mathbb{R}^{n})$ we define:
@@
    If $S_{V,\psi,\rho,\varphi}>0$ on $\mathbb{R}_{+}$ and $(\varphi^{N}r^{n})^{\prime}\neq 0$ a.e on $\mathbb{R}_{+}$, then $\varphi$ is a solution to (6.2) if and only if the function $\tau$ satisfies
@@ 6.2. Existence of solutions with freely specified mean curvature
Although the Euclidean space is the most familiar setting, existence results for (1.2) are almost restricted to the CMC case. This is because, apart from the limit equation criterion — which, as explained at the beginning of this section, has not yet been established for the conformal equations on AF manifolds — there are essentially only two approaches available for proving existence of solutions in the non-CMC regime. The first relies on the smallness of $\frac{d\tau}{\tau}$ to prevent blow-up of solutions, corresponding to the near-CMC case. The second is based on the smallness of $\sigma$ and $\rho$, leading to small solutions and is commonly called the small TT-tensor method. Unfortunately, despite the results in [14], none of these methods is convincingly applicable on $\mathbb{R}^{n}$. On the one hand, the weighted Poincaré inequality does not allow us to assume that $\frac{d\tau}{\tau}$ is too small; on the other hand, the fact that solutions converges to $1$ at infinity breaks the smallness assumption on solutions required to implement the small TT-tensor method.
@@
with $\varphi\longrightarrow 1$ at infinity. The close analytical similarity of these equations to the vacuum (1.2) and their importance in providing a global perspective to the solvability of (1.2) have been explained briefly in the previous sections. For that, similarly to what we have done previously, to establish convincing evidence supporting that the conformal method is still an useful tool for constructing solutions to the constraint equations (1.1) on AF manifolds, we study (6.4) particularly in the simple model where $(M,g)$ is Euclidean and $(\tau,\rho)$ are assumed to be radial. In this context, motivated by the proof of Theorem 4.10, the main result we obtain in this section is that as long as $\rho$ is $C^{1}$-regular, the equations (6.4) are always solvable, with solutions uniformly bounded independently of $\tau$. Before presenting the precis statement and providing the proof, we first introduce the associated Banach spaces on $(\mathbb{R}^{n},\sigma_{\text{Euc}})$ as follows. Given an integer $l\geq 0$, a Hölder exponent $\alpha\in[0,1]$, and a decay exponent $\beta>0$, we will use the weighted Hölder spaces $C^{l,\alpha}_{-\beta}$ to capture asymptotic of functions and tensors near infinity. For $\alpha=0$, we will write $C^{l}_{-\beta}$ instead of $C^{l,0}_{-\beta}$. The weighted norm convention we are using is that the $C^{s,\alpha}_{-\beta}$ norm is given by
@@
In view of [13, Section 5], we obtain that $\widetilde{T}$ is continuous compact and it is clear that a fixed point of $\widetilde{T}(1,)$ is a solution to (6.4). Now, we set
@@
$\frac{\rho^{2}(\varphi^{N}r^{n})^{\prime}}{\varphi^{2N}}=-(\rho^{2}r^{2n})\bigg(\frac{1}{\varphi^{N}r^{n}}\bigg)^{\prime},$
@@
6.3. The sharpness of the rate of decay in the positive mass theorem